OpenStudy (anonymous):

Help please What mass, in grams, of oxygen gas (O2) is contained in a 8.5 liter tank at 32.0 degrees Celsius and 3.23 atmosphere?

OpenStudy (anonymous):

@dan815 @aaronq @Luigi0210

OpenStudy (anonymous):

Use pV= mRT/Mr, where in is in Pascals, V in m^3, R is 8.31, T is temperature in Kelvin , m is required mass and Mr is molecular mass. Think you can solve it now?

OpenStudy (anonymous):

yea i got that part but i need to know the steps

OpenStudy (thadyoung):

Make sure you're using the proper gas constant. I would use 0.0821 instead since you're already given liters and atm. Just convert temperature to absolute temp (which is kelvin)

OpenStudy (anonymous):

this is what i got so far (3.23 atm)(8.5l)=(n)(0.082atm*L/mole*K)(274.15k)

OpenStudy (anonymous):

and im stuck @thadyoung @Hoslos

OpenStudy (thadyoung):

You need to rearrange like you would in algebra, and solve for \(n\) which is moles then you can use that to calculate grams. Since you know that oxygen has a molar mass of 16, which is given in grams per mole

OpenStudy (thadyoung):

Rearrange: \(\sf PV = nRT \Rightarrow \color{red}{\frac{PV}{RT}}= n\)

OpenStudy (anonymous):

okay so (3.23)(8.5L)/(0.082atm*L/mole*K)(274.15)?

OpenStudy (anonymous):

@thadyoung

OpenStudy (thadyoung):

Close, but where did you come up with 274.15 for your temperature?

OpenStudy (anonymous):

i converted the celcius into kelvin

OpenStudy (anonymous):

so 32.0C is 274.15 K

OpenStudy (thadyoung):

32 degree's celcius is \(not\) 274.15 K Do you know the conversion?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

its 305.15 @thadyoung

OpenStudy (anonymous):

so its 3.23 atm x 8.5L / (0.08206 Latm/molK x305K)

OpenStudy (anonymous):

am i correct?

OpenStudy (thadyoung):

Yes.

OpenStudy (anonymous):

now whats next?

OpenStudy (anonymous):

this is where i get stuck

OpenStudy (thadyoung):

Well, you have moles. Remember from algebra, if you have this: \(\sf \large \frac{a}{b} \times \frac{b}{z} = \frac{a}{z} \) because the two b's cancel out? \(\sf \large \frac{a}{\cancel{b}} \times \frac{\cancel{b}}{z} = \frac{a}{z} \)

OpenStudy (anonymous):

yes

OpenStudy (thadyoung):

It is the same thing with units. You have moles of Oxygen (O\(_2\)) And when you see your periodic table, you know that the units are in molar mass. Which menas you have n moles of O\(_2\) \(\times\) \(\sf \frac{15.99~g}{mol}\)

OpenStudy (thadyoung):

That's the mass for oxygen.

OpenStudy (thadyoung):

Now, does that make sense?

OpenStudy (anonymous):

ohh k yea kind of

OpenStudy (thadyoung):

Notice that you have units mole that can cancel out?

OpenStudy (thadyoung):

This is still simple algebra.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so the two moles cancel out?

OpenStudy (thadyoung):

Yes.

OpenStudy (anonymous):

so how would the formula look now

OpenStudy (thadyoung):

exactly how i wrote it up there. try it.

OpenStudy (anonymous):

okay so one mole of ox weighs 32g so i multiply that by what 15.99g?

OpenStudy (anonymous):

@thadyoung

OpenStudy (anonymous):

im so confused? :(

OpenStudy (anonymous):

whats ar ethe number of moles?

OpenStudy (thadyoung):

Put it in your calculator! I'm not doing your work for you. I already showed you how to do it.